OpenStudy (anonymous):

A box contains 3 green marbles and 2 blue marbles. 2 marbles are chosen without replacement. If both are blue, you win $10. If just one is blue you lose $3. Otherwise you lose $1. What is your expected gain or loss?

OpenStudy (campbell_st):

2/7*1/6 *10 + 1/7*5/6*-3 + 5/7*4/6*-1

OpenStudy (anonymous):

how did you get that?

OpenStudy (campbell_st):

probability of selection 1 blue of 1st pick 2/7 blue on 2nd pick. 1/6 so its 2/7*1/6*10 next its 1st pick probability blue = 2/7 2nd pick not blue = 5/6 oops forgot probability not blue 1st pick 5/7 probability blue 2nd pick 2/6 and 1st pick P not blue 5/7 2nd pick not blue = 4/6 the the expextation is the probability times financial return 2/7*1/6 *10 + 2/7*5*6*-3 + 5/7*2/6*-3 + 5/7*4/6*1

OpenStudy (anonymous):

how do you get the 7 for your denominator? There are only 5 marbles.

OpenStudy (campbell_st):

there are 7 marbles... in the box

OpenStudy (campbell_st):

if a selection is made and not put back in the box... there is only 6 to select from

OpenStudy (anonymous):

thanks!